# Vaccination in Metropolitan Areas: Should the Capital Be Prioritized in Vaccine Allocation?

### v1

@@ 3.3 Performance of time-variable vaccination strategies
In this section, we perform numerical simulations that incorporate vaccination as a time-variable control function. By integrating vaccination strategies into our metropolitan area model, the goal is to provide a comprehensive understanding of disease dynamics and control measures. We start by examining the interaction between two cities, as shown in Figure 11. This figure illustrates the optimal trajectory and control for a capital city with a population $10$ times greater than the second city and a higher infection rate. Vaccination begins on the $100$th day following the arrival of the first infected individual in the capital. Notably, during the initial three weeks, the optimal solution allocates all vaccines to the capital, administering them at the maximum possible rate according to the weekly cap. Subsequently, as the susceptible population in the capital decreases to around $60\%$, it becomes advantageous to start vaccinating the second city, which has a lower daily cap. The fourth graph illustrates the bang-bang [33] behavior of the effective vaccination rate $u_{i}S_{i},$ this is, in which $u_{i}^{*}(t)S_{i}^{*}(t)\in\{0,v^{\text{max}}_{i}\}$ holds for $i=1,2$, along the whole interval $[0,T].$ The gap of the weekly cap constraint is depicted in Figure 12. We assign a high unity cost to hospitalization $c_{h}=1000$ in contrast to a relatively low unity cost for vaccination, with $c_{v}=0.01$ to reflect the burden of hospitalization. Both healthcare resources and patient well-being are comparatively way more expensive than the minor inconvenience and price of vaccination. Furthermore, the benefits of vaccination are amplified when administered on a large scale. This simulation took around $33$ seconds to run.
Figure 11: Optimal trajectories and control for two-city interaction: this figure shows the optimal trajectory of the proportion of susceptible and infectious individuals in two cities, the capital in orange and the second city in blue. The fourth subplot illustrates the variables $u_{i}(t)S_{i}(t)\in\{0,v^{\text{max}}_{i}\},\,i=1,2$. The settings for this experiment are: $\beta=(0.25,0.18)$, $\alpha=0.64$, $p_{21}=0.2$, $n_{1}=10^{6}$ and $n_{2}=10^{5}$. For the optimal control problem, $c_{v}=0.01$, $c_{h}=1000$, $v^{\mathrm{max}}_{1}=v^{\mathrm{max}}_{2}=0.8/42$ and a weekly cap allowing the vaccination of at most $1/13$ of the population per week.
Figure 12: Weekly constraint analysis: this figure illustrates the function $D(t)-\sum_{k=1}^{K}n_{k}V_{k}(t)$ for the experiment presented in Figure 11. In red, we fix the value 00 to give a reference.
@@
Figure 13: Optimal trajectories and control for two-city interaction for uniform infection rates: this figure shows the optimal trajectory of the proportion of susceptible and infectious individuals in two cities and the optimal control using the same settings as Figure 11 except for $\beta=(0.25,0.25)$.
@@
Finally, we consider four different scenarios: first infection in the second city (Scenario (I)), uniform initial conditions for the control problem (Scenario (II)), and two that express a higher transmissibility in the capital, but a lower vaccination rate limitation (Scenarios (III) and (IV)). The result is shown in Figure 14. It establishes that the vaccination always starts in the capital in all scenarios, even with the infection starting in the second city. Moreover, Scenarios (III) and (IV) confirm that this behavior does not come from the maximum effective vaccination rate, highlighting that even though we can vaccinate more rapidly in the second city, the capital starts first.
@@
Figure 14: Optimal effective vaccination rate for two-city interaction: this figure shows four different scenarios. Scenario (I) considers the same setting as Figure 11 except for the first infection happening in the second city and $v_{1}^{\max}=0.9/42$ (while $v_{2}^{\max}=0.8/42$). Scenario (II) considers $I_{1}(0)=I_{2}(0)=0.05$ and $R_{1}(0)=R_{2}(0)=0.02$. Scenarios (III) and (IV) fix the same parameters of Figure 11 but with different maximum rates $v_{1}^{\max}$ and $v_{2}^{\max}$. In (III), we set $v_{1}^{\mathrm{max}}=0.8/42<v_{2}^{\mathrm{max}}=0.9/42$, while in (IV), $v_{1}^{\mathrm{max}}=0.6/42<v_{2}^{\mathrm{max}}=0.9/42$.
@@
Having considered the simplified two-city case, we extend our analysis to more cities in the metropolitan area. We first observe that the dimensionality of the problem escalates quickly, which makes efficient implementations a challenging aspect yet to be addressed in the literature. Lemaitre et al. (2022[17] considers a similar optimal control problem and offers a viable solution, which we could further explore in future research.
@@
For an simulation with $K=5$ cities, refer to Figure 15. The results contained in the image are parallel to those in the two-city scenario in which the vaccines are preferentially allocated to the capital in contrast to other cities, vaccinating as fast as possible saturating the constraints. During the fourth week, when the susceptible population is reduced to less than $60\%$ in the capital, the vaccination expands to the second city, which has the second highest transmission rate. The behavior of the control function aligns again with the bang-bang type of solution. In total, the experiment required approximately $3$ minutes and $31$ iterations.
@@
Figure 15: Optimal trajectories and control functions for $5$ cities: this figure presents the optimal trajectories and controls for a problem involving $5$ cities. The settings mirror those in 21(a) in the Appendix, with the following specifications: $c_{v}=0.01$, $c_{h}=1000$, and $r_{h}=0.1$. The model allows for a weekly vaccination of up to $1/20$ of the susceptible population. The maximum rate of vaccinated individuals is set at $v_{1}^{\max}=\dots=v_{5}^{\max}=0.8/42$.
@@
Figure 16: Comparison between optimal and constant vaccination strategies for different values of $\beta_{1}$: this bar plot figure compares the total number of infections in the metropolitan area, which is calculated by integrating the new infections and taking the mean over the cities, for two scenarios. The first, in blue, solves the optimal control problem (9), and the second, in orange, chooses the best constant vaccination rate across all cities.
@@ 3.4 Performance of feedback practical solution
$\frac{d}{dt}(V_{i}+S_{i}+I_{i})=\mu R_{i}-\gamma I_{i}\leq 0\implies V_{i}(t)+S_{i}(t)\leq V_{i}(t)+S_{i}(t)+I_{i}(t)\leq V_{i}(0)+S_{i}(0)+I_{i}(0)\leq 1$
@@
Figure 17: Comparison between optimal and practical solution in the metropolitan area: this figure presents the optimal trajectories and effective vaccination rate for the metropolitan area considering the settings of Figure 15. It compares the optimal solution with a more practicable solution, which is only based on the knowledge of vaccinated individuals. We observe that the effective vaccination rate here is the daily proportion of vaccinated individuals in the metropolitan area as a whole.
@@ 3.5 Evaluating Vaccination Strategies in the Rio de Janeiro Metropolitan Area
Figure 18: Comparative Analysis of Optimal Vaccination Strategies in the Rio de Janeiro Metropolitan Area: this figure illustrates the optimal vaccination strategies for Rio de Janeiro, through both the original and approximated $P$ matrices. The settings are $\alpha=0.64$, a maximum vaccination rate of $0.6/56$ per city, and weekly vaccine shipments for vaccinating up to $1/20$ of the total susceptible population. The transmission rate $\beta$ is randomly chosen between 00 and $0.3$, and sorted by city population size. Vaccination starts when the susceptible population falls below 95% following the initial infection in the capital. The approximation of $P$ in a metropolitan context is calculated by normalizing $p_{i1}$ and $p_{ii}$ by its sum.
@@
The parameter $\beta$ is proportional to the average contact rate of a random person. It is generally an increasing function of the population density. Given the heterogeneous density in Rio de Janeiro’s metropolitan area, which includes large green spaces like the Tijuca Forest in the capital city, it would be not enough to only consider $\beta$ as proportional to the average density of each city. For this reason, we consider random values for $\beta$ and sort them by population size.
Vaccination in the capital and the other cities is compared in Figure 19. This analysis tracks on a daily basis the number of cities that vaccinate while the capital is not, i.e., $u_{1}(t)=0$ and some $u_{i}(t)\neq 0$ for $i\geq 2$. We notice that most of the time, either the capital is vaccinating or none of the cities are. The first deviation from this fact happens when the susceptible population is less than $60\%$, which corroborates with our findings on the behavior of the optimal control function.
Figure 19: Analysis of Effective Vaccination Rates in the Rio de Janeiro Metropolitan Area: this figure presents two graphs – the red graph represents the proportion of susceptible individuals in Rio de Janeiro city, while the gray bars show the number of cities vaccinating when the capital is not. The settings of this experiment are identical to those in Figure 18.
@@
Figure 20: Disease Progression in the Rio de Janeiro Metropolitan Area under Various Vaccination Scenarios: This figure displays the cumulative proportion of infections via a heatmap, utilizing the framework established in Figure 18. Data sourced from Brazilian Institute of Geography and Statistics (IBGE).
@@ Appendix A The basic reproduction number for general compartmental models
where $F$ is non-negative, $V$ is a non-singular $M$-matrix [^2] A square matrix $B$ is an $M$-matrix if it can be expressed as $B=sI-P$, where $P$ is a matrix with non-negative elements, and $s$ is a real scalar such that $s\geq\rho(P)$. If $s>\rho(P)$, then $B$ is a non-singular $M$-matrix. If $s=\rho(P)$, it is a singular $M$-matrix. and the eigenvalues of $J_{4}$ have positive real part [39, Lemma 1]. The matrix $FV^{-1}$ is named the next generation matrix for the model and
@@ A.1 Computing the basic reproduction number for our model
The unique disease-free equilibrium of system (1) sets $S_{i}=1$ and $I_{i}=R_{i}=0$ for all cities $i=1,\dots,K$, which simplifies $\mathcal{S}$ to the identity matrix. In the modified model (5) with a constant vaccination rate in the population, the dynamics of compartments $I_{i}$ are unchanged, which leads to the same calculation of the basic reproduction number. However, the disease-free equilibrium is redefined as
@@ Appendix B Proofs
is positively invariant under the flow of system (1).
@@ C.2 Epidemic values for different structures of the transition matrix
In the context of multiple cities, we consider a scenario of $K=5$ cities. We first examine five distinct mobility scenarios: (I) metropolitan area where $10\%$ of the residents of each city work in the capital; (II) metropolitan area where $40\%$ of each city works in the capital; (III) all cities are interconnected, where 60% of the residents work in their city, while the remaining are distributed evenly across the other four cities; (IV) a variant of a metropolitan area where 30% of each city’s residents work in the capital and 10% of the capital’s workforce is employed in the other cities; (V) no commuting between cities. Our findings, summarized in Table C.2
@@ C.4 Impact of vaccination in the peak size
Although $\mathcal{R}_{0}$ provides a good understanding of the impact of vaccination on epidemic dynamics, it only informs about the asymptotic behavior of the dynamics. For a finite time horizon, analyzing the peak size of the epidemic provides crucial insights into the optimal vaccination strategy. This is the goal of 21(a), which presents four different scenarios under a uniform vaccination policy. We have chosen $s=42$ days for this experiment. The peak size of the baseline case, when no vaccine is introduced, is about 21% of the population. To reduce it to around 15%, as shown in orange, over 60% of the population needs to be vaccinated. Early vaccination can significantly reduce the peak size of the epidemic, as illustrated in 21(b), which shows the impact of the vaccination starting date on the peak size.

### v2

@@ 3.3 Performance of time-variable vaccination strategies
In this section, we perform numerical simulations that incorporate vaccination as a time-variable control function. By integrating vaccination strategies into our metropolitan area model, the goal is to provide a comprehensive understanding of disease dynamics and control measures. We start by examining the interaction between two cities, as shown in Figure 11. This figure illustrates the optimal trajectory and control for a capital city with a population $10$ times greater than the second city and a higher infection rate. Vaccination begins on the $100$th day following the arrival of the first infected individual in the capital.
Figure 11: Optimal trajectories and control for two-city interaction.
Optimal trajectory of the proportion of susceptible and infectious individuals in two cities, the capital in orange and the second city in blue. The fourth subplot illustrates the variables $u_{i}(t)S_{i}(t)\in\{0,v^{\text{max}}_{i}\},\,i=1,2$. The settings for this experiment are: $\beta=(0.25,0.18)$, $\alpha=0.64$, $p_{21}=0.2$, $n_{1}=10^{6}$ and $n_{2}=10^{5}$. For the optimal control problem, $c_{v}=0.01$, $c_{h}=1000$, $v^{\mathrm{max}}_{1}=v^{\mathrm{max}}_{2}=0.8/42$ and a weekly cap allowing the vaccination of at most $1/13$ of the population per week.
Notably, during the initial two weeks, the optimal solution allocates all vaccines to the capital, administering them at the maximum possible rate according to the weekly cap. Subsequently, as the susceptible population in the capital decreases to around $60\%$, it becomes advantageous to start vaccinating the second city, which has a lower daily cap. The fourth graph illustrates the bang-bang [33] behavior of the effective vaccination rate $u_{i}S_{i},$ this is, in which $u_{i}^{*}(t)S_{i}^{*}(t)\in\{0,v^{\text{max}}_{i}\}$ holds for $i=1,2$, along the whole interval $[0,T].$ The gap of the weekly cap constraint is depicted in Figure 12. We assign a high unity cost to hospitalization $c_{h}=1000$ in contrast to a relatively low unity cost for vaccination, with $c_{v}=0.01$ to reflect the burden of hospitalization. Both healthcare resources and patient well-being are comparatively way more expensive than the minor inconvenience and price of vaccination. Furthermore, the benefits of vaccination are amplified when administered on a large scale. This simulation took around $33$ seconds to run.
Figure 12: Weekly constraint analysis.
Function $D(t)-\sum_{k=1}^{K}n_{k}V_{k}(t)$ for the experiment presented in Figure 11.
@@
Figure 13: Optimal trajectories and control for two-city interaction for uniform infection rates.
Optimal trajectory of the proportion of susceptible and infectious individuals in two cities and the optimal control using the same settings as Figure 11 except $\beta=(0.25,0.25)$.
@@
Finally, we consider four different scenarios: first infection in the second city (Scenario (I)), uniform initial conditions for the control problem (Scenario (II)), and two that express higher transmissibility in the capital, but a lower vaccination rate limitation (Scenarios (III) and (IV)). The result is shown in Figure 14. It establishes that the vaccination always starts in the capital in all scenarios, even with the infection starting in the second city. Moreover, Scenarios (III) and (IV) confirm that this behavior does not come from the maximum effective vaccination rate, highlighting that even though we can vaccinate more rapidly in the second city, the capital starts first.
Figure 14: Optimal effective vaccination rate for two-city interaction.
@@
Scenario (I) considers the same setting as Figure 11 except for the first infection happening in the second city and $v_{1}^{\max}=0.9/42$ (while $v_{2}^{\max}=0.8/42$). Scenario (II) considers $I_{1}(0)=I_{2}(0)=0.05$ and $R_{1}(0)=R_{2}(0)=0.02$. Scenarios (III) and (IV) fix the same parameters of Figure 11 but with different maximum rates $v_{1}^{\max}$ and $v_{2}^{\max}$. In (III), we set $v_{1}^{\mathrm{max}}=0.8/42<v_{2}^{\mathrm{max}}=0.9/42$, while in (IV), $v_{1}^{\mathrm{max}}=0.6/42<v_{2}^{\mathrm{max}}=0.9/42$.
@@
Having considered the simplified two-city case, we extend our analysis to more cities in the metropolitan area. We first observe that the dimensionality of the problem escalates quickly, which makes efficient implementations a challenging aspect yet to be addressed in the literature. Lemaitre et al. (2022[17] consider a similar optimal control problem and offer a viable solution, which we could further explore in future research.
@@
For a simulation with $K=5$ cities, refer to Figure 15. The results contained in the image are parallel to those in the two-city scenario in which the vaccines are preferentially allocated to the capital in contrast to other cities, vaccinating as fast as possible saturating the constraints. During the fourth week, when the susceptible population is reduced to less than $60\%$ in the capital, the vaccination expands to the second city, which has the second highest transmission rate. The behavior of the control function aligns again with the bang-bang type of solution. In total, the experiment required approximately $3$ minutes and $31$ iterations.
Figure 15: Optimal trajectories and control functions for $5$ cities.
@@
Optimal trajectories and controls for a problem involving $5$ cities. The settings mirror those in Figure C.22, with the following specifications: $c_{v}=0.01$, $c_{h}=1000$, and $r_{h}=0.1$. The model allows for a weekly vaccination of up to $1/20$ of the susceptible population. The maximum rate of vaccinated individuals is set at $v_{1}^{\max}=\dots=v_{5}^{\max}=0.8/42$.
@@
Figure 16: Comparison between optimal and constant vaccination strategies for different values of $\beta_{1}$.
Comparison among the total number of infections in the metropolitan area, which is calculated by integrating the new infections and taking the mean over the cities, for two scenarios. The first, in blue, solves the optimal control problem (9), and the second, in orange, chooses the best constant vaccination rate across all cities.
@@ 3.4 Performance of feedback practical solution
$\frac{d}{dt}(V_{i}+S_{i}+I_{i})=\mu R_{i}-\gamma I_{i}\leq 0,$
which implies
$V_{i}(t)+S_{i}(t)\leq V_{i}(t)+S_{i}(t)+I_{i}(t)\leq V_{i}(0)+S_{i}(0)+I_{i}(0)\leq 1$
@@
Figure 17: Comparison between optimal and practical solution in the metropolitan area.
Optimal trajectories and effective vaccination rate for the metropolitan area considering the settings of Figure 15. It compares the optimal solution with a more practicable solution, which is only based on the knowledge of vaccinated individuals. We observe that the effective vaccination rate here is the daily proportion of vaccinated individuals in the metropolitan area as a whole.
@@ 3.5 Evaluating Vaccination Strategies in the Rio de Janeiro Metropolitan Area
Figure 18: Comparative Analysis of Optimal Vaccination Strategies in the Rio de Janeiro Metropolitan Area.
Optimal vaccination strategies for Rio de Janeiro, through both the original and approximated $P$ matrices. The settings are $\alpha=0.64$, a maximum vaccination rate of $0.6/56$ per city, and weekly vaccine shipments for vaccinating up to $1/20$ of the total susceptible population. The transmission rate $\beta$ is randomly chosen between 00 and $0.3$, and sorted by city population size. Vaccination starts when the susceptible population falls below 95% following the initial infection in the capital. The approximation of $P$ in a metropolitan context is calculated by normalizing $p_{i1}$ and $p_{ii}$ by its sum.
@@
The parameter $\beta$ is proportional to the average contact rate of a random person. It is generally an increasing function of the population density. Given the heterogeneous density in Rio de Janeiro’s metropolitan area, which includes large green spaces like the Tijuca Forest in the capital city, it would not be enough only to consider $\beta$ as proportional to the average density of each city. For this reason, we consider random values for $\beta$ and sort them by population size.
Vaccination in the capital and the other cities is compared in Figure 19. This analysis tracks on a daily basis the number of cities that vaccinate while the capital is not, i.e., $u_{1}(t)=0$ and some $u_{i}(t)\neq 0$ for $i\geq 2$. We notice that most of the time, either the capital is vaccinating or none of the cities are. The first deviation from this fact happens when the susceptible population is less than $60\%$, which corroborates our findings on the behavior of the optimal control function.
Figure 19: Analysis of Effective Vaccination Rates in the Rio de Janeiro Metropolitan Area.
The red graph represents the proportion of susceptible individuals in Rio de Janeiro city, while the gray bars show the number of cities vaccinating when the capital is not. The settings of this experiment are identical to those in Figure 18.
@@
Figure 20: Disease Progression in the Rio de Janeiro Metropolitan Area under Various Vaccination Scenarios.
Cumulative proportion of infections via a heatmap, with the framework established in Figure 18. Data sourced from Brazilian Institute of Geography and Statistics (IBGE).
@@ Appendix A The basic reproduction number for general compartmental models
where $F$ is non-negative, $V$ is a non-singular $M$-matrix and the eigenvalues of $J_{4}$ have positive real part [39, Lemma 1]. A square matrix $B$ is an $M$-matrix if it can be expressed as $B=sI-P$, where $P$ is a matrix with non-negative elements, and $s$ is a real scalar such that $s\geq\rho(P)$. If $s>\rho(P)$, then $B$ is a non-singular $M$-matrix. If $s=\rho(P)$, it is a singular $M$-matrix. The matrix $FV^{-1}$ is named the next generation matrix for the model and
@@ A.1 Computing the basic reproduction number for our model
The unique disease-free equilibrium of system (LABEL:eq:robot-dance-model) sets $S_{i}=1$ and $I_{i}=R_{i}=0$ for all cities $i=1,\dots,K$, which simplifies $\mathcal{S}$ to the identity matrix. In the modified model (5) with a constant vaccination rate in the population, the dynamics of compartments $I_{i}$ are unchanged, which leads to the same calculation of the basic reproduction number. However, the disease-free equilibrium is redefined as
@@ Appendix B Proofs
is positively invariant under the flow of system (LABEL:eq:robot-dance-model).
@@ C.2 Epidemic values for different structures of the transition matrix
In the context of multiple cities, we consider a scenario of $K=5$ cities. We first examine five distinct mobility scenarios: (I) metropolitan area where $10\%$ of the residents of each city work in the capital; (II) metropolitan area where $40\%$ of each city works in the capital; (III) all cities are interconnected, where 60% of the residents work in their city, while the remaining are distributed evenly across the other four cities; (IV) a variant of a metropolitan area where 30% of each city’s residents work in the capital and 10% of the capital’s workforce is employed in the other cities; (V) no commuting between cities. Our findings are summarized in Table C.2.
@@ C.4 Impact of vaccination in the peak size
Although $\mathcal{R}_{0}$ provides a good understanding of the impact of vaccination on epidemic dynamics, it only informs about the asymptotic behavior of the dynamics. For a finite time horizon, analyzing the peak size of the epidemic provides crucial insights into the optimal vaccination strategy. This is the goal of 22(a), which presents four different scenarios under a uniform vaccination policy. We have chosen $s=42$ days for this experiment. The peak size of the baseline case, when no vaccine is introduced, is about 21% of the population. To reduce it to around 15%, as shown in orange, over 60% of the population needs to be vaccinated. Early vaccination can significantly reduce the peak size of the epidemic, as illustrated in 22(b), which shows the impact of the vaccination starting date on the peak size.